### v1

# Mpemba effect on non-equilibrium active Markov chains

Abstract
We study the Mpemba effect on a non-equilibrium Markov chain that mimics the run and tumble motion of an active particle in a discrete energy landscape. The broken detailed balance, rendered by the activity, gives rise to a unique anomalous relaxation in the system which is distinctly different than the typical equilibrium systems. We observe that the activity can both suppress or induce the Mpemba effect. Furthermore, we report an oscillatory Mpemba effect where the relaxation trajectories, emanating from the hot and cold initial conditions, cross each other multiple times and this occurs due to the emergence of complex eigenvalues in the relaxation spectrum due to the activity. Our work reveals a possible pathway for studying the Mpemba effect in active living systems where broken detailed balance is crucial to achieve many biological functions.
The Mpemba effect refers to the anomalous relaxation phenomena where an initially hotter system equilibrates faster than an initially cooler system, when both systems are quenched to the same low temperature [1]. Although the phenomena was first observed in the case of water namely, hot water freezing faster [1, 2, 3, 4, 5, 6, 7, 8, 9] than an equal volume of warm water exposed to the same condition, it is now a well-observed phenomena that has been probed through both theoretical and experimental means across various systems [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 24, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. A particular frontier that remains notably unexplored is the effect of activity in the anomalous relaxation dynamics (besides the work by Schwarzendahl and Löwen on active colloids [29]). Activity is a fingerprint of non-equilibrium systems which unlike their equilibrium counterpart violate detailed balance and this results in many interesting features such as non-Boltzmannian steady states, clustering, spontaneous aggregation, motility induced phase transition in active self-propelled systems, sense and adaption for high fidelity DNA transcription, structural organization in cells, folding of macromolecules [53]. In this paper, we aim to understand how the broken detailed balance, specifically induced due to the activity, can manifest in the anomalous relaxation dynamics such as the Mpemba effect.
Figure 1: Illustration of an active Markov chain model capturing the dynamics of a run and tumble motion of an active walker in a discrete 3-state energy landscape. Panel (a): The hopping rates $R_{ij}$ that satisfy detailed balance and in effect, characterize the rates of diffusion of an active walker among the energy states. Panel (b) depicts the corresponding energy landscape where $B_{ij}$’s are the energy barriers between states with energies $E_{i}$ and $E_{j}$. Panel (c): The activity of the walker is characterised by the rates $\gamma$ and $\epsilon$, where $\gamma$ is the flipping rate between the states $\uparrow$ and $\downarrow$ – equivalent to the persistence between the two states, and $\epsilon$ is the state-dependent rate of unidirectional hopping that resembles the velocity of the active walker. In essence, $\gamma\neq 0$ is the non-equilibrium active component that drives the system away from equilibrium.
In particular, we show via a rigorous analysis that an active system can display a rich behaviour in the relaxation dynamics such as the emergence of oscillations in the transient & the activity -induced or -suppressed Mpemba effect. The results are intriguing as the presence of such behaviours in the eigenspectrum or in the relaxation dynamics has no counterpart in regular equilibrium setups. Although a similar oscillatory behaviour in the relaxation has been reported recently in a two-level quantum system where such behaviour is attributed to the role played by the presence of exceptional points and complex eigenvalues [54], the presence of such behaviours in the classical models is rather unique which we explore herein. Finally, the presence of non-equilibrium steady states and the oscillatory relaxations raises plausible questions on the adaptable measures (pertaining only to the truncated eigenspectrum) that are often used to quantify the Mpemba effect in equilibrium systems [27, 22].
Model and framework: Consider a setup of a discrete three state active Markov chain (Fig. 1). The intrinsic hopping rate of a random walker from the state $j$ to $i$ is denoted by $R_{ij}$ and vice-versa. These rates satisfy detailed balance and are determined at a fixed bath temperature $T_{b}$ by the following relation
$\begin{split}R_{ij}=e^{-\frac{B_{ij}-Ej}{k_{B}T_{b}}},\text{ for }i\neq j,\\ R_{ji}=-\sum_{k\neq i}R_{ki},\text{ for }i=j,\end{split}$ (1)
where $E_{i}$ are the energy levels of the three states and $B_{ij}$ are the barrier heights between any two states $i$ and $j$. These rates usually characterize the rates of diffusion of the walker among the different energy states. To introduce the activity, we assume that there is an internal spin degree of freedom (denoted by $\uparrow$ and $\downarrow$) associated with each state. The spin can flip with a rate $\gamma$ which characterizes the persistence of the walker to be in that particular internal state. Depending on its current internal state ($\uparrow$ or $\downarrow$), the walker hops either to the clockwise or the anti-clockwise site at a rate $\epsilon$, thus characterizing the velocity of the active walker [55]. Hence, this model mimics a prototypical run and tumble motion of an active walker [55] on a discrete energy landscape.
The central quantity required to understand the anomalous relaxation is the joint occupation probability $p_{i,\sigma}(t)$ which estimates the probability for the walker to be in the site $i$ in $\sigma$ state, where $\sigma=\{\uparrow,\downarrow\}$. The master equations for $p_{i,\sigma}(t)$ at time $t$ can be constructed from the knowledge of the hopping and active rates as
$\frac{dp_{1,\sigma}}{dt}=$ $(R_{12}\pm\epsilon)p_{2,\sigma}+(R_{13}\mp\epsilon)p_{3,\sigma}-(R_{21}+R_{31})p_{1,\sigma}$
$-\gamma p_{1,\sigma}+\gamma p_{1,\bar{\sigma}},$ (2)
and similarly, for the sites $i=2,3$. Here, $\bar{\sigma}$ denotes the conjugate of $\sigma$. To solve for the occupation probability, it is useful to define the following quantities
$P_{i}=p_{i,\uparrow}+p_{i,\downarrow},$
$P^{*}_{i}=p_{i,\uparrow}-p_{i,\downarrow},$ (3)
where $P_{i}$ denotes the total probability of being at site $i$ summed over the internal spin states $\uparrow$ or $\downarrow$, and $P^{*}_{i}$ can be understood as the polarisation at site $i$ and its positive value denotes the preference of state $\uparrow$ over $\downarrow$, and vice-versa. Equation (3) can now be written in a concise matrix form
$\frac{d\boldsymbol{P}(T,t)}{dt}=\boldsymbol{W}(T_{b})\boldsymbol{P}(T,t),$ (4)
where the vector $\boldsymbol{P}=(P_{1},P_{2},P_{3},P^{*}_{1},P^{*}_{2},P^{*}_{3})^{\mathbb{T}}$ denotes the instantaneous probability distributions of the system at time $t$ and $\boldsymbol{W}(T_{b})$ is the transition matrix that is determined at the bath temperature $T_{b}$. Here, the initial condition is chosen to be a steady state probability distribution $\boldsymbol{P}(T,t=0)$ characterized by a different bath temperature $T\neq T_{b}$ to which the system is initially prepared before being quenched to the final bath temperature $T_{b}$. If the transitions between microscopic states are pairwise-balanced (like in Eq. (1)), the system reaches a thermodynamic equilibrium precluding net flux among the states. On the other hand, if the transitions are not balanced, the system can exhibit flux loops, and moreover attains a non-equilibrium steady state.
Figure 2: Eigenvalues $\{\lambda\}$-s of the transition matrix $\boldsymbol{W}$ of an active Markov chain (with broken detailed balance): Presence of both real and complex eigenvalues is observed as we change the activity parameters. The transition from real to mixed eigenvalues as $\gamma$ is varied for fixed $\epsilon=1.0$ [(in panel (a)] and $\epsilon=9.0$ [(in panel (b)].
Evidently, with the above-mentioned dynamics with induced activity (hence, broken detailed balance), the time evolution of the distribution $\boldsymbol{P}(T,t)$ attains a steady state distribution, $\text{lim}_{t\to\infty}\boldsymbol{P}(T,t)=\boldsymbol{\pi}(T_{b})$ at large times, corresponding to the final quenched temperature of the bath $T_{b}$. During the relaxation process, the time evolution of the initial distribution $\boldsymbol{P}(T,t=0)$ is obtained by solving the time-dependent Eq. (4) and is given by
$\boldsymbol{P}(T,t)=\boldsymbol{\pi}(T_{b})+\sum^{6}_{i=2}a_{i}(T,T_{b})\boldsymbol{v}_{i}e^{\lambda_{i}t},$ (5)
where $\lambda_{i}$ and $\boldsymbol{v}_{i}$ are the eigenvalues and right eigenvectors of the transition matrix $\boldsymbol{W}(T_{b})$ respectively such that
$\boldsymbol{W}(T_{b})\boldsymbol{v}_{i}=\lambda_{i}\boldsymbol{v}_{i}.$ (6)
The eigenvalues follow the order $\lambda_{1}=0>\text{Re}\lambda_{2}\geq\text{Re}\lambda_{3}\ldots,$ where $\lambda_{1}=0$ is the largest eigenvalue that corresponds to the eigenvector $\boldsymbol{\pi}(T_{b})$. All other eigenvalues have negative real parts. The coefficient $a_{i}(T,T_{b})$ in Eq. (5) is the overlap coefficient of the $i^{th}$ eigenvector with the initial distribution $\boldsymbol{P}(T,t=0)$.
If the transition rates satisfy detailed balance, it is guaranteed that all the eigenvalues are real [56]. On the other hand, for non-equilibrium dynamics with broken detailed balance, the eigenspecturm can be mixed with the presence of both real and complex eigenvalues. Here too, we observe that the eigenvalues turn out to be complex as one varies the activity parameter $\gamma$. Figure 2 illustrates such a transition in the eigenspectrum as a result of the change in the active parameters $\epsilon$ and $\gamma$.
Figure 3: Activity induced and suppressed Mpemba effect. Panel (a): The Mpemba effect is absent in the passive case ($\epsilon=\gamma=0$) while the presence of activity ($\epsilon=2$ and $\gamma=4$) induces the Mpemba effect (shown in inset). The model parameters are: $E_{1}=0.0$, $E_{2}=0.1$, $E_{3}=0.3$, $B_{12}=B_{21}=0.4$, $B_{13}=B_{31}=0.5$, $B_{23}=B_{32}=0.4$, $T_{h}=4.0$, $T_{c}=1.0$ and $T_{b}=0.1$. Panel (b) shows activity suppressed Mpemba effect (shown in inset) while it is present in the passive case. Here, the active rates are $\epsilon=1$ and $\gamma=2$. The model parameters in this case are: $E_{1}=0.2$, $E_{2}=0.3$, $E_{3}=0.5$, $B_{12}=B_{21}=0.9$, $B_{13}=B_{31}=0.6$, $B_{23}=B_{32}=0.9$, $T_{h}=3.0$, $T_{c}=1.0$ and $T_{b}=0.1$. Although the measure $L_{1}$-norm has been used as a reference, the results are found to be independent of the distance measures [SM].
The Mpemba effect: To interpret the Mpemba effect, let us consider two systems: first one $A$, initially in a steady state corresponding to a hot temperature $T_{h}$ and second one $B$, initially in a different steady state corresponding to cold temperature $T_{c}$ where $T_{h}>T_{c}$. These initial steady states are denoted by $\boldsymbol{\pi}(T_{h})$ and $\boldsymbol{\pi}(T_{c})$ respectively. Now imagine that both $A$ and $B$ are quenched at once to a common bath temperature, $T_{b}$, where $T_{h}>T_{c}>T_{b}$. Eventually, both of them will relax to the common distribution $\boldsymbol{\pi}(T_{b})$ given long enough time. The Mpemba effect is said to exist if $A$ reaches the final steady state faster than $B$ during the transient/relaxation process. This relaxation can be quantified in terms of the distance function $D[\boldsymbol{\mathcal{P}}(t),\boldsymbol{\Pi}(T_{b})]$ which measures the instantaneous distance of the state-wise occupation probabilities $\boldsymbol{\mathcal{P}}(t)=(P_{1},P_{2},P_{3})^{\mathbb{T}}$ from their respective final steady state distributions $\boldsymbol{\Pi}(T_{b})=(\pi_{1},\pi_{2},\pi_{3})^{\mathbb{T}}$. There are many well-adapted measures that exist in the literature namely the norm $L_{1}(t)=\sum_{i=1}^{3}|P_{i}(t)-\pi_{i}(T_{b})|$ and Kullback-Leibler (KL) divergence $KL(t)=\sum_{i=1}^{3}P_{i}(t)\ln(P_{i}(t)/\pi_{i})$. Thus, if one has $D[\boldsymbol{\Pi}(T_{h}),\boldsymbol{\Pi}(T_{b})]>D[\boldsymbol{\Pi}(T_{c}),\boldsymbol{\Pi}(T_{b})]$ initially for $T_{h}>T_{c}$ followed by $D[\boldsymbol{\Pi}(T_{h}),\boldsymbol{\Pi}(T_{b})]<D[\boldsymbol{\Pi}(T_{c}),\boldsymbol{\Pi}(T_{b})]$ at a later time, we infer that the Mpemba effect exists.
Activity harnessed Mpemba effect: To highlight the effect of activity, we first set $\epsilon=\gamma=0$ and consider the passive discrete three-state Markov chain [27]. The identification of the Mpemba effect in this case can be made from the behaviour of the coefficient $|a_{2}(T,T_{b})|$ [as in Eq. (5)] [27]. In particular, we choose a set of parameters of the system (namely $R_{ij}$-s) where the Mpemba effect is absent. With the present parameters as the activity is introduced, we find an emergence of the Mpemba effect [see Fig. 3]. We also note a converse effect where the introduction of activity suppresses an already existing Mpemba effect [see Fig. 3 for more details]. Notably, this behavior remains quite robust under other distance measures [see SM].
Figure 4: Illustration of the oscillatory Mpemba effect across different distance measures: (a) the $L_{1}$-norm, and (b) the Kullback-Leibler (KL) divergence for a system with broken detailed balance. The complex eigenspecturm can lead to oscillatory relaxations and distance measures. Here, the model parameters are: $E_{1}=0.0$, $E_{2}=0.1$, $E_{3}=0.3$, $B_{12}=B_{21}=0.4$, $B_{13}=B_{31}=0.5$, $B_{23}=B_{32}=0.4$, $\gamma=8$, $\epsilon=9$ whereas the temperatures of the hot, cold and the bath temperatures are chosen to be: $T_{h}=5.0$, $T_{c}=0.5$ and $T_{b}=0.1$, respectively.
The qualitative reasoning of the occurrence of the Mpemba effect in the energy landscape with multiple local minima is given in terms of the effective trapping of the initially prepared colder system in the metastable states leading to its slower relaxation. In the context of the current model, a passive random walker initiated at a hot temperature does not experience any trap and explores the energy landscape much faster than the initially prepared colder system. On the contrary, an active random walker, be it initiated at a hot or cold temperature or even for a significant barrier height of the landscape, may not experience any metastable state for intermediate values of its activity or persistence as it can now explore the phase space with additional frequency characterized by the active rates. Such a phenomena suppresses the Mpemba effect in the presence of activity. However, the activity can also act otherwise. For example, the active random walker can experience momentary trap in the energy landscape due its persistence in a particular state leading to an activity induced Mpemba relaxation. We note that a similar phenomena was also observed in active colloidal systems diffusing in a potential landscape [29]. Naturally, the activity induced or suppressed behavior is a complex convolution of the initial temperature of the particle, barrier heights of the underlying energy landscape and the active, diffusive rates.
Figure 5: Oscillatory Mpemba effect characterised by the zero crossings in $\Delta L_{1}(t):=L^{hot}_{1}(t)-L^{cold}_{1}(t)$ for two initially prepared hot and cold systems as they relax to a common steady state. We measure $\Delta L_{1}(t)$ as a function of time for varying flipping rate $\gamma$ which shows a pronounced effect of the activity. The oscillations tend to go down with increased $\gamma$ as the active walker becomes a diffusive one in this limit [55]. Here, we have set $\epsilon=9.0$ throughout the analysis.
Oscillatory Mpemba effect: As shown in Fig. (2), the eigenvalues of the transition matrix $\boldsymbol{W}(T_{b})$ can be complex. As a result, both the time evolution of the probability distribution and the distance measures will have oscillatory terms. Figure 4 illustrates the time evolution of two different distance measures, namely $L_{1}$ and KL, for two systems $A$ and $B$ that are initiated at two different temperatures $T_{h}$ and $T_{c}$ respectively with $T_{h}>T_{c}$. We show that unlike the previous results of the Mpemba effect where the distance measures decay monotonically with time, here the distance measures show non-monotonic oscillatory-like decay across all the measures. As a result, the trajectories of the initially hot and cold system cross multiple times. This also indicates that the so-called $|a_{2}(T,T_{b})|$-criterion which is used for monotonic relaxations can not be a rational measure for non-equilibrium systems with complex eigenvalues.
To have a better representation for the oscillations, we plot the difference of the $L_{1}$-norm for the hot and the cold system namely $\Delta L_{1}(t):=L^{hot}_{1}(t)-L^{cold}_{1}(t)$ as a function of time $t$ for different activity rates $\gamma$ keeping $\epsilon$ fixed as shown in Fig. 5. The plots shows that $\Delta L_{1}$ crosses the origin multiple times before it asymptotically relaxes to zero. The number of oscillations decreases as function of $\gamma$ for a fixed $\epsilon$ and thus, for the higher values of the activity $\gamma$, we can observe a reentrant diffusive like behavior in the relaxation of the trajectories. This is not surprising as in the large flipping rate $\gamma$ limit, the active walker behaves like a passive one as can also be seen from the effective diffusion constant [55, 57].
Figure 6: Multiple crossings in the relaxation trajectories can emerge even with real eigenspectrum of the active Markov chain (with $\gamma=\epsilon=1$). The result is independent of distance measures used as shown here with: (a) $L_{1}$-norm, and (b) Kullback-Leibler (KL) divergence. The results shown are for model parameters: $E_{1}=0.0$, $E_{2}=0.1$, $E_{3}=0.3$, $B_{12}=B_{21}=0.6$, $B_{13}=B_{31}=0.1$, $B_{23}=B_{32}=0.6$, whereas the temperatures of the hot, cold and the bath temperatures are chosen to be: $T_{h}=10.0$, $T_{c}=1.0$ and $T_{b}=0.1$, respectively.
On the contrary, with real eigenspectrum, is it always ensured to have a single crossing of trajectories? We consider one such case but find that the distance functions can cross twice before they asymptotically merge with each other as illustrated in Fig. 6. Since there are even crossings, the hotter system lags behind the initially colder system leading to no net Mpemba effect in the large time. The same conclusion is drawn from the measure $|a_{2}(T,T_{b})|$ as well [see Fig. S2 in the SM]. Note that such a scenario is also observed in the non-equilibrium model of driven granular gases [58, 49] (albeit not universally but depending on the choice of the distance measure). Generically, this behavior can be attributed to the preparation of the systems in non-equilibrium conditions as is the case for the active systems here.
Conclusions: In this paper, we have investigated the role of broken detailed balance in the anomalous relaxation phenomena. We unveil that the broken detailed balance conferred by the activity in a Markov chain system can showcase a myriad of unique relaxation phenomena such as the activity -induced and -suppressed Mpemba effect, and oscillatory Mpemba effect where relaxation trajectories cross multiple times. The latter is owing to the emergence of the complex eigenspectrum of the transition matrix in the presence of activity.
Qualitatively, some of these phenomena can be reasoned with a heuristic description of ‘activity induced exploration on the energy landscape’. For instance, a cold walker can overcome metastabilities, kinetic traps or bottlenecks with the aid of active rates suppressing the Mpemba effect. However, activity can also act in favour of the Mpemba effect by trapping the walker of the initially colder system in a particular state with strong persistence. This would lead to the activity induced Mpemba effect. An intermediate persistence renders the walker to switch its state intermittently and thus, one would observe a non-monotonic relaxation behavior. Further analysis can shed light on this physical picture.
While it is known that the active fluctuations can lead to an extensive variety of chemical and biological functions in living systems, its direct role in the quenching & relaxation dynamics pertaining to the Mpemba effect is yet to be characterized. Drawing inspiration from this simple analysis of an active Markov chain, we posit that some of these ideas maybe relevant to coarse-grained living systems such as fluctuating enzymes [59], proteins, or molecular machines [60] which continuously dissipate energy with the surrounding or exchange energy with the other molecules [53].
It would be interesting to see whether experimental setups can be designed to capture such Mpemba like non-equilibrium phenomena from the transport & relaxation of bio-molecules through the engineered structures and cell biological systems [61] or even in simple active colloidal systems [62]. We believe that a synergy between the theory and experiment can provide a deeper conceptual understanding of this non-equilibrium phenomena.
Acknowledgements: AP acknowledges research support from the DST SERB Start-up Research Grant Number SRG/2022/000080 and the Department of Atomic Energy, Government of India. AP thanks Amit Kumar Chatterjee and Hisao Hayakawa for useful discussions and for sharing relevant references on the quantum Mpemba effect. AP also thanks Udo Seifert for insightful discussions.